# Croppable Knowledge Graph Embedding

### v1

@@ 4.2 Evolutionary Improvement Mechanism
The hard (ground-truth) label is the other important supervision signal during training in knowledge distillation DBLP:journals/corr/HintonVD15 . High-dimensional sub-models need to master triples that low-dimensional sub-models can not learn well, that is, high-dimensional sub-models need to correctly predict those positive (negative) triples that are wrongly predicted to be negative (positive) by low-dimensional sub-models. In MED, for a given triple $(h,r,t)$, the optimization weight in sub-model $M_{i}$ for it depends on the triple score output by the previous sub-model $M_{i-1}$.
For a positive triple, the optimization weight of the model $M_{i}$ for it is negatively correlated with its score by the model $M_{i-1}$. Specifically, the higher its score from the model $M_{i-1}$ (meaning that $M_{i-1}$ has been able to correctly judge it as a positive sample), the lower the optimization weight of the model $M_{i}$ for it, and the lower its score from the model $M_{i-1}$ (meaning that $M_{i-1}$ wrongly judges it as a negative sample), the higher the optimization weight of the model $M_{i}$ for it because $M_{i-1}$ cannot predict this triple well. The optimization weight of $M_{i}$ for the positive triple is
$pos_{h,r,t}^{i}=\frac{\exp w_{1}/s_{(h,r,t)}^{i-1}}{\sum_{(h,r,t)\in T_{batch}}\exp w_{1}/s_{(h,r,t)}^{i-1}}\ \texttt{if}\ 1<i\leqslant n\ ;\quad\frac{1}{|T_{batch}|}\ \texttt{if}\ i=1,$ (2)
where $s_{(h,r,t)}^{i-1}$ is the score for triple $(h,r,t)$ output by the sub-model $M_{i-1}$, $T_{batch}$ is the set of positive triples within a batch, and $w_{1}$ is a learnable scaling parameter. Conversely, for a negative triple, the optimization weight of the model $M_{i}$ for it is positively correlated with its score by the model $M_{i-1}$. The optimization weight of $M_{i}$ for the negative triple is
$neg_{h,r,t}^{i}=\frac{\exp w_{2}\cdot s_{(h,r,t)}^{i-1}}{\sum_{(h,r,t)\in T_{batch}^{-}}\exp w_{2}\cdot s_{(h,r,t)}^{i-1}}\ \texttt{if}\ 1<i\leqslant n\ ;\quad\frac{1}{|T_{batch}^{-}|}\ \texttt{if}\ i=1,$ (3)
where $T_{batch}^{-}$ is the set of negative triples within a batch, and $w_{2}$ is a learnable scaling parameter.
Therefore, the evolutionary improvement loss of the sub-model $M_{i}$ is
$L_{EI}^{i}=-\sum_{(h,r,t)\in\mathcal{T}\cup\mathcal{T}^{-}}pos_{h,r,t}^{i}\cdot y\log\sigma(s_{(h,r,t)}^{i})+neg_{h,r,t}^{i}\cdot(1-y)\log(1-\sigma(s_{(h,r,t)}^{i})),$ (4)
where $\sigma$ is the Sigmoid activation function, $y$ is the ground-truth label of the triple $(h,r,t)$, and it is $1$ for positive triples and 00 for negative ones. In each sub-model, different hard (ground-truth) label loss weights are set for different triples, and the high-dimensional sub-model will pay more attention to learn the triple that the low-dimensional sub-model can not learn well.
@@ 4.3 Dynamic Loss Weight
Since MED involves the optimization of multiple sub-models, we set dynamic loss weights during training. Initially, low-dimensional sub-models prioritize learning from high-dimensional sub-models to improve performance. This means low-dimensional sub-models rely more on soft label information, so for low-dimensional sub-models, evolutionary improvement loss should account for less than mutual learning loss. Conversely, high-dimensional sub-models should focus more on capturing knowledge that low-dimensional models lack, while mitigating the impact of low-quality outputs from low-dimensional models to maintain their good performance, that is, high-dimensional sub-models rely more on hard label information. So for high-dimensional sub-models, evolutionary improvement loss should account for more than mutual learning loss. For a teacher-student pair, their mutual learning loss acts on both teacher and student models simultaneously, so the effect of mutual learning loss for them is theoretically the same. We set different evolutionary improvement loss weights for different sub-models, and the final training loss function of MED is
@@ 5 Experiment
We evaluate MED on typical KGE and GLUE benchmarks and particularly answer the following research questions: (RQ1) Is it capable for MED to train a croppable KGE at once that multiple sub-models of different dimensions can be cropped from it and all achieve promising performance? (RQ2) Can MED finally achieve parameter-efficient KGE models? (RQ3) Does MED work in real-world applications? (RQ4) Can MED be extended to other neural networks besides KGE?
@@ 5.1.1 Dataset and KGE methods
MED is universal and can be applied to any KGE method with a triple score function, we select three commonly used KGE methods as examples: TransE DBLP:conf/nips/BordesUGWY13 , RotatE DBLP:conf/iclr/SunDNT19 and PairRE DBLP:conf/acl/ChaoHWC20 , the triple score functions are described in Table 1.
Table 2: Statistics of datasets.
@@
We conduct comparison experiments on two common KG completion benchmark datasets WN18RR DBLP:conf/emnlp/ToutanovaCPPCG15 and FB15K237 DBLP:conf/aaai/DettmersMS018 and two more larger-scale KGs CoDEx-L DBLP:conf/emnlp/SafaviK20 and YAGO3-10 DBLP:conf/cidr/MahdisoltaniBS15 . Besides, we apply our MED on a real-world large-scale e-commerce social knowledge graph (SKG) involving more than 50 million triples of social records by about 7 million users in the Taobao platform in real application scenarios. Table 2 shows the statistics of the datasets.
@@ 5.1.2 Evaluation Metric
For the link prediction task, we adopt standard metrics MRR and Hit@$k$ $(k=1,3,10)$ in the filtered setting DBLP:conf/nips/BordesUGWY13 . We use Effi DBLP:conf/aaai/ChenZYZGPC23 , that is MRR/#P (#P is the number of parameters), to quantify the parameter efficiency of models. We use f1-score and accuracy for the user labeling task, and normalized discounted cumulative gain ndcg@$k(k=5,10)$ for the product recommendation task.
@@ 5.1.3 Implementation
For the link prediction task, we set $d_{n}=640$ for the highest-dimensional sub-model $M_{n}$ and $d_{1}=10$ for the lowest-dimensional sub-model $M_{1}$. We set $n=64$ and the same dimension size gap $10$ for every pair of neighbor sub-models so that there are a total of $64$ available sub-models of different dimensions from 10 to 640 in our croppable KGE model. The dimension of sub-model $M_{i}(i=1,2...,64)$ is $10\times i$. For the user labeling and product recommendation task, we set $n=3$ and train the croppable KGE containing 3 sub-models: $M_{1}$ with $d_{1}=10$ for mobile phone (MB) terminals that are limited by storage and computing resources, $M_{2}$ with $d_{2}=100$ for the personal computer (PC), and $M_{3}$ with $d_{3}=500$ for the platform’s servers. We initialize the learnable scaling parameters $w_{i},w_{2}$ and $w_{3}$ in (2), (3) and (5) to 1. We implement MED by extending OpenKE DBLP:conf/emnlp/HanCLLLSL18 , an open-source KGE framework based on PyTorch. We set the batch size to $1024$ and the maximum training epoch to $3000$ with early stopping. For each positive triple, we generate $64$ negative triples by randomly replacing its head or tail entity with another entity. We use Adam DBLP:journals/corr/KingmaB14 optimizer with a linear decay learning rate scheduler and perform a search on the initial learning rate in $\{0.0001,0.0005,0.001,0.01\}$. We train all sub-models simultaneously by optimizing the uniformly sampled sub-models from the full Croppable model in each step.
@@ 5.1.4 Baselines
For each required dimension $d_{r}$, we extract the first $d_{r}$ dimensions from our croppable KGE as the target model and compare it to the KGE models obtained by 7 baselines of the following 3 types:
  • Directly training the target KGE model of requirement dimension $d_{r}$, referred to as 1) DT. The directly trained highest-dimensional KGE model ($d_{r}=d_{n}$) is marked as $M_{max}^{DT}$.
  • Extracting the first $d_{r}$ dimensions from $M_{max}^{DT}$ as the target model, referred to as 2) Ext. Besides, we update $M_{max}^{DT}$ by assessing the importance of each one of 640 dimensions and arranging them in descending order before extracting as DBLP:conf/iclr/MolchanovTKAK17 ; DBLP:conf/acl/VoitaTMST19 : 3) Ext-L, the importance for each dimension of $M_{max}^{DT}$ is the variation of KGE loss on validation set after removing it; and 4) Ext-V, the importance for each dimension is the average absolute of its parameter weights of all entities and all relations.
  • Distilling the target KGE by KD methods: 5) BKD DBLP:journals/corr/HintonVD15 is the most basic one by minimizing the KL divergence of the output distributions of the teacher and student; 6) TA DBLP:conf/aaai/MirzadehFLLMG20 uses a medium-size teaching assistant (TA) model as a bridge for size gap, where TA model has the same dimension as the directly trained one whose MRR is closest to the average MRR of the teacher and student; and 7) DualDE DBLP:conf/wsdm/ZhuZCCC0C22 compresses KGE by optimizing the teacher and student simultaneously. We do not compare with MulDE DBLP:conf/www/Wang0MS21 , which uses multiple low-dimensional different KGE models as teachers to aggregate the knowledge of different KGE models into one rather than compress a high-dimensional KGE. In these baselines, $M_{max}^{DT}$ is the teacher, and other settings including hyperparameters are the same as their original papers.
@@ 5.2 Performance Comparison
MED outperforms baselines in almost all settings, especially for the extremely low dimensions. On WN18RR with $d$=10, MED achieves an improvement of 14.9% and 15.1% on TransE, 8.4% and 6.6% on RotatE, 29.4% and 10.6% on PairRE compared with the best MRR and Hit@10 of baselines. We can observe a similar phenomenon on FB15K237. This benefits from the rich knowledge sources of low-dimensional models in MED: For sub-model $M_{i}$, $M_{i+1}$ is the teacher directly next to it, while $M_{i+2}$ can also indirectly affect $M_{i}$ by directly affecting $M_{i+1}$. Theoretically, all higher-dimensional sub-models can finally transfer their knowledge to low-dimensional sub-models through stepwise propagation. Although such stepwise propagation may have negative effects on high-dimensional models by bringing low-quality knowledge from low-dimensional sub-models, the evolutionary improvement mechanism in MED weakens the damage and makes high-dimensional ones still
achieve competitive performance than directly trained KGEs as in Fig. 3. We also find that Ext-based methods perform extremely unstable: Ext, Ext-L, and Ext-V work worse than DT except on WN18RR with TransE, indicating that only considering the importance of each dimension is not enough to guarantee the performance of all sub-models. More results and ablation studies are in Appendix A and Appendix B.
5.3 Parameter efficiency of MED
In Table 4, we compare our sub-models of suitable low dimensions to parameter-efficient KGEs especially proposed for large-scale KGs including NodePiece DBLP:conf/iclr/0001DWH22 and EARL DBLP:conf/aaai/ChenZYZGPC23 . In the case that the number of model parameters is roughly equivalent, the performance of the sub-models of MED exceeds that of the specialized parameter-efficient KGE methods. This demonstrates sub-models of our method are parameter efficient. More importantly, it can provide parameter-efficient models of different size for applications.
@@ 5.3 Parameter efficiency of MED
5.4 MED in real applications
We apply the trained croppable KGE with TransE on SKG to three real applications: the user labeling task on servers and the product recommendation task on PCs and mobile phones. Table 5 shows that our croppable user embeddings substantially exceed all baselines including directly trained (DT), the best baseline DualDE, and a common dimension reduction method in industry principal components
Table 5: Results on SKG.
@@ 5.4 MED in real applications
Method acc. f1 ndcg@5 ndcg@10 ndcg@5 ndcg@10
DT 0.889 0.874 0.411 0.441 0.344 0.361
PCA - - 0.417 0.447 0.392 0.418
DualDE - - 0.423 0.456 0.404 0.433
MED 0.893 0.879 0.431 0.465 0.422 0.451
analysis (PCA) on $M^{DT}_{max}$. Notably, the excellent performance on the mobile phone task (which can only carry embeddings with a maximum dimension of 10 limited by storage and computing resources) demonstrates the enormous practical value of our approach. More application details are in Appendix C.
5.5 Extend MED to Neural Networks
To verify the extensibility of our method to other neural networks, we take the language model BERT DBLP:conf/naacl/DevlinCLT19 as an example. To ensure the consistency of the experimental environment as much as possible, we uniformly adopt distillation methods implemented based on Hugging Face Transformers DBLP:conf/emnlp/WolfDSCDMCRLFDS20 as baselines. Following previous works DBLP:conf/emnlp/SunCGL19 ; DBLP:journals/corr/abs-1903-12136 ; DBLP:journals/eswa/JungKNK23 ; DBLP:conf/acl/ZhouXM22 , we do not use pre-training distillation settings and only distill at the fine-tuning stage. More experimental details are in Appendix D.
Table 6: Results on the dev set of GLUE. The results of knowledge distillation methods for BERT4 and BERT6 are reported by DBLP:journals/eswa/JungKNK23 ; DBLP:conf/acl/ZhouXM22 and the results reported by us.
@@ 5.5 Extend MED to Neural Networks
Table 6 shows the results on the development set of GLUE DBLP:conf/iclr/WangSMHLB19 . We compare MED with other KD models under similar speedup or a comparable number of parameters. The results show that MED achieves competitive performance on most tasks compared to BERT-specialized KD methods. In addition, when compared to HAT DBLP:conf/acl/WangWLCZGH20 , which shares the most similar model architecture to ours, sub-models of MED outperform HAT across three different parameter quantities. Specifically, sub-models with 54M, 17.5M, and 6.36M parameters achieve average $16.3\%$, $21.7\%$ and $19.7\%$ improvements respectively.

### v2

@@ 4.2 Evolutionary Improvement Mechanism
In knowledge distillation Hinton et al. (2015), the hard (ground-truth) label is a key supervision signal during training. High-dimensional sub-models need to master triples that low-dimensional sub-models can not, that is correctly predicting positive (negative) triples mispredicted as negative (positive) by low-dimensional sub-models. In MED, for a triple $(h,r,t)$, sub-model $M_{i}$’s optimization weight for it depends on triple score from $M_{i-1}$.
For a positive triple, $M_{i}$’s optimization weight is negatively correlated with the score from $M_{i-1}$. If $M_{i-1}$ gives a high score (correctly identifying it as positive), $M_{i}$’s optimization weight for it is low. If $M_{i-1}$ gives a low score (misidentifying it as negative), $M_{i}$’s optimization weight is high as $M_{i-1}$ can’t predict it well. The optimization weight of $M_{i}$ for the positive triple is
$\begin{split}pos_{h,r,t}^{i}=\left\{\begin{array}[]{ll}1/|T_{batch}|,&i=1\\ \frac{\exp w_{1}/s_{(h,r,t)}^{i-1}}{\sum\limits_{(h,r,t)\in\mathcal{T}_{b}}\exp w_{1}/s_{(h,r,t)}^{i-1}},&1<i\leqslant n,\end{array}\right.\end{split}$ (3)
where $s_{(h,r,t)}^{i-1}$ is the score for triple $(h,r,t)$ from $M_{i-1}$, $\mathcal{T}_{b}$ is positive triple set in a batch, and $w_{1}$ is a learnable scaling parameter. Conversely, for a negative triple, $M_{i}$’s optimization weight is positively correlated with the score from $M_{i-1}$, as also used in  Sun et al. (2019b). The optimization weight of $M_{i}$ for negative triple is
$\begin{split}neg_{h,r,t}^{i}=\left\{\begin{array}[]{ll}1/|T_{batch}^{-}|,&i=1\\ \frac{\exp w_{2}\cdot s_{(h,r,t)}^{i-1}}{\sum\limits_{(h,r,t)\in\mathcal{T}_{b}^{-}}\exp w_{2}\cdot s_{(h,r,t)}^{i-1}},&1<i\leqslant n,\end{array}\right.\end{split}$ (4)
where $\mathcal{T}_{b}^{-}$ is negative triple set in a batch, and $w_{2}$ is a learnable scaling parameter. The evolutionary improvement loss of the sub-model $M_{i}$ is
$L_{EI}^{i}=-\sum_{(h,r,t)\in\mathcal{T}\cup\mathcal{T}^{-}}pos_{h,r,t}^{i}\cdot y\log\sigma(s_{(h,r,t)}^{i})$ (5)
$+neg_{h,r,t}^{i}\cdot(1-y)\log(1-\sigma(s_{(h,r,t)}^{i})),$
In each sub-model, different hard-label loss weights are set for different triples, and high-dimensional sub-models focus more on triples low-dimensional ones can’t learn well.
@@ 4.3 Dynamic Loss Weight
As MED optimizes multiple sub-models, we use dynamic loss weights during training. At first, low-dimensional sub-models prioritize learning from high-dimensional ones to improve performance, relying more on soft label information. Thus, for them, evolutionary improvement loss should be less than mutual learning loss. Conversely, high-dimensional sub-models should focus more on what low-dimensional ones can’t correctly predict and reduce the impact of low-quality low-dimensional outputs. They rely more on hard label information, so their evolutionary improvement loss should exceed mutual learning loss. For a teacher-student pair, the mutual learning loss affects both models equally. We set different evolutionary improvement loss weights for sub-models, and the final training loss function of MED is
@@ 5 Experiment
We evaluate MED on typical KGE and GLUE benchmarks and address these research questions: (RQ1) Can MED train a croppable KGE at once, with multiple differently-dimensional sub-models cropped from it, all performing well? (RQ2) Can MED yield parameter-efficient KGE models? (RQ3) Does MED work in real-world applications? (RQ4) Can MED be extended to non-KGE neural networks?
@@ 5.1.1 Dataset and KGE methods
MED is universal and applicable to any KGE method with a triple score function. We take 4 common KGE methods in Table 1 as examples: TransE Bordes et al. (2013), SimplE Kazemi and Poole (2018), RotatE Sun et al. (2019b) and PairRE Chao et al. (2021).
@@
Table 2: Statistics of datasets.
We conduct comparison experiments on on four KG completion benchmark datasets: two common ones, WN18RR Toutanova et al. (2015) and FB15K237 Dettmers et al. (2018), and two larger-scale ones, CoDEx-L Safavi and Koutra (2020) and YAGO3-10 Mahdisoltani et al. (2015). In real-world scenarios, we apply MED to a large-scale e-commerce social knowledge graph (SKG) from Taobao, which has over 50 million social record triples from about 7 million users. Table 2 presents dataset statistics.
@@ 5.1.2 Evaluation Metric
For link prediction task, we adopt standard metrics MRR and Hit@$k$ $(k=1,3,10)$ with filter setting Bordes et al. (2013). For a test triple $(h,r,t)$, we construct candidate triples by replacing $h$ with all entities and calculate the triple score rank of $(h,r,t)$ among candidate triples as the head prediction rank $rank_{h}$. Likewise, we get the tail prediction rank $rank_{t}$. We average $rank_{h}$ and $rank_{t}$ as $(h,r,t)$’s final rank. MRR is the mean reciprocal rank of all test triples, Hit@$k$ is the percentage of test triples with rank $\leq k$. We use Effi=MRR/#P (#P is the parameter number) Chen et al. (2023) to measure models’ parameter efficiency. We use f1-score and accuracy for user labeling task, and normalized discounted cumulative gain ndcg@$k(k=5,10)$ for product recommendation task.
@@ 5.1.3 Implementation
For the link prediction task, we set $d_{n}=640$ for the highest-dimensional sub-model $M_{n}$ and $d_{1}=10$ for the lowest-dimensional sub-model $M_{1}$. We set $n=64$ and the dimension gap $10$ for every pair of neighbor sub-models. There are a total of $64$ available sub-models of different dimensions from 10 to 640 in our croppable KGE model. The dimension of $M_{i}(i=1,2...,64)$ is $10\times i$. For the user labeling and product recommendation task, we set $n=3$ and train the croppable KGE containing 3 sub-models: $M_{1}$ with $d_{1}=10$ for mobile phone (MB) terminals that are limited by storage and computing resources, $M_{2}$ with $d_{2}=100$ for the personal computer (PC), and $M_{3}$ with $d_{3}=500$ for the platform’s servers. We initialize the learnable scaling parameters $w_{1},w_{2}$ and $w_{3}$ in (3), (4) and (6) to 1. We implement MED by extending OpenKE Han et al. (2018), an open-source KGE framework based on PyTorch. We set the batch size to $1024$ and the maximum training epoch to $3000$ with early stopping. For each positive triple, we generate $64$ negative triples by randomly replacing its head or tail entity with another entity. We use Adam Kingma and Ba (2015) optimizer with a linear decay learning rate scheduler and perform a search on the initial learning rate in $\{0.0001,0.0005,0.001,0.01\}$. We train all sub-models simultaneously by optimizing the uniformly sampled sub-models from the full Croppable model in each step.
@@ 5.1.4 Baselines
For each required dimension $d_{r}$, we extract the first $d_{r}$ dimensions from our croppable KGE as the target model and compare it to the KGE models obtained by 8 baselines of the following 3 types:
  • Directly training the target KGE model of requirement dimension $d_{r}$, referred to as 1) DT. The directly trained highest-dimensional KGE model ($d_{r}=d_{n}$) is denoted as $M_{max}^{DT}$.
  • Extracting the first $d_{r}$ dimensions from $M_{max}^{DT}$ as the target model, called 2) Ext. We update $M_{max}^{DT}$ by arranging 640 dimensions in descending order based on their importance before extracting as Molchanov et al. (2017); Voita et al. (2019): 3) Ext-L, the importance for each dimension of $M_{max}^{DT}$ is the variation of KGE loss on validation set after removing it; and 4) Ext-V, the importance for each dimension is the average absolute of parameter weights of all entities and relations.
  • Distilling the target KGE by KD methods: 5) BKD Hinton et al. (2015) is the most basic one by minimizing the KL divergence of the output distributions of teacher and student; 6) TA Mirzadeh et al. (2020) uses a medium-size teaching assistant (TA) model as a bridge for size gap, where TA model has the same dimension as the directly trained one whose MRR is closest to the average MRR of teacher and student. We also compare with two KD methods proposed for KGE, which have similar configurations to ours, i.e. compressing high-dimensional teacher into low-dimensional student: 7) DualDE Zhu et al. (2022) considers the mutual influences between teacher and student and optimizes them simultaneously; 8) IterDE Liu et al. (2023) enables the KGE model to alternately act as student and teacher so that knowledge can be transferred smoothly between high-dimensional teacher and low-dimensional student. In these baselines, $M_{max}^{DT}$ is the teacher, and other settings including hyperparameters are the same as their original papers.
@@ 5.2 Performance Comparison
MED outperforms baselines in most settings, especially at extremely low dimensions. On WN18RR with $d$=10, MED achieves an improvement of 14.9% and 15.1% on TransE, 8.4% and 6.6% on RotatE, 29.4% and 10.6% on PairRE than the best MRR and Hit@10 of baselines. A similar trend is seen on FB15K237. This success stems from MED’s rich knowledge sources for low-dimensional models: For sub-model $M_{i}$, $M_{i+1}$ is a direct teacher, and $M_{i+2}$ can indirectly influence $M_{i}$ via $M_{i+1}$. In theory, all higher-dimensional sub-models can transfer knowledge to lower-dimensional ones through stepwise propagation. Although this propagation may harm high-dimensional models with low-quality knowledge from low-dimensional ones, MED’s evolutionary improvement mechanism mitigates the damage, allowing high-dimensional models to compete with directly trained KGEs (Fig. 3). We also note that Ext-based methods are highly unstable, performing worse than DT in most cases, suggesting that dimension importance alone doesn’t ensure sub-model performance.
5.3 Parameter efficiency of MED
In Table 4, we compare our sub-models of suitable low dimensions to parameter-efficient KGEs especially proposed for large-scale KGs including NodePiece Galkin et al. (2022) and EARL Chen et al. (2023). With a similar number of model parameters, MED’s sub-models outperform specialized parameter-efficient KGE methods, showing their parameter efficiency. More crucially, it can offer different-sized parameter-efficient models for applications.
@@ 5.3 Parameter efficiency of MED
Method train time acc. f1 ndcg@5 ndcg@10 ndcg@5 ndcg@10
DT 103h 0.889 0.874 0.411 0.441 0.344 0.361
PCA - - - 0.417 0.447 0.392 0.418
DualDE 195h - - 0.423 0.456 0.404 0.433
MED 53h 0.893 0.879 0.431 0.465 0.422 0.451
Table 5: Results on SKG.
5.4 MED in real applications
We apply the trained croppable KGE with TransE on SKG to three real-world applications: user labeling on servers and product recommendation on PCs and mobile phones. Table 5 shows that our croppable user embeddings outperform all baselines, including directly trained (DT), the best baseline DualDE, and a common dimension reduction method in industry principal components analysis (PCA) on $M^{DT}_{max}$. Notably, the strong performance in the mobile phone task (constrained by storage and computing resources to a max embedding dimension of 10) highlights the great practical value of MED. More application details are in Appendix C.
@@ 5.4 MED in real applications
Table 6: Results on the dev set of GLUE. The results of knowledge distillation methods for BERT4 and BERT6 are reported by Jung et al. (2023); Zhou et al. (2022a) and the results reported by us.
5.5 Extend MED to Neural Networks
To test our method’s extensibility to other neural networks, we use BERT Devlin et al. (2019) as an example. We use distillation methods based on Hugging Face Transformers Wolf et al. (2020) as baselines. As in previous worksSun et al. (2019a); Tang et al. (2019); Jung et al. (2023); Zhou et al. (2022a), we perform distillation at the fine-tuning stage. See Appendix D for more details. Table 6 shows the results on the GLUE development setWang et al. (2019). We compare MED with other KD models of similar speedup or parameter number. MED performs competitively on most tasks against BERT-specialized KD methods. Compared to HAT Wang et al. (2020a), which has a similar architecture to ours, MED’s sub-models outperform HAT across three parameter levels. Sub-models with 54M, 17.5M, and 6.36M parameters show average improvements of $16.3\%$, $21.7\%$ and $19.7\%$ respectively.